if sec0=3 and tan0=2sqr root 2 what are the values of
cot0
sin0
cos^2 0

Answer by nabla(475)

By the use of 0 you mean theta,θ, correct?
(1) sec(θ)=3=1/cos(θ)
(2) tan(θ)=2sqrt(2)=1/cot(θ)=sin(θ)/cos(θ)

a. cot(θ)=?
From (2) we notice that cot(θ)=1/(2sqrt(2))=sqrt(2)/4
b. sin(θ)=?
From (1) and (2) we see that sin(θ)=cos(θ)tan(θ)=(1/3)*2sqrt(2)=(2/3)sqrt(2)
c. cos^2(θ)=?
From (1) we can see that cos(θ)=1/3, thus cos^2(θ)=1/9